Apply Cardano's method for finding the roots of 2x³+3x²-8x-12=0
1
Expert's answer
2019-07-08T10:16:23-0400

"2x^3+3x^2-8x-12=0""x^3+(3\/2)x^2-4x-6=0""(ax^3+bx^2+cx+d=0)"

Substitute

"x=y-(b\/3)=y-(1\/2)"

"(y-(1\/2))^3+3\/2(y-(1\/2))^2-4(y-1\/2)-6=0"

"y^3-(19\/4)y-(15\/4)=0 (y^3+py+q=0)"

Formula Cardano:

"Q=(p\/3)^3+(q\/2)^2, A=((-q\/2)+Q^(1\/2))^(1\/3), B=((-q\/2)-Q^(1\/2))^(1\/3)"

"y1=A+B, y2=(-(A+B)\/2)+3^(1\/2)(A-B\/2)i, y3=(-(A+B)\/2)-3^(1\/2)(A-B\/2)i"

"Q=-5959\/64"

Since the value is less than 0, 3 complex numbers will be solutions of the equation


"A=(15\/8+(5959^(1\/2)\/8)i)^1\/2""B=(15\/8-(5959^(1\/2)\/8)i)^1\/2""y1=(15\/8-(5959^(1\/2)\/8)i)^1\/2+(15\/8+(5959^(1\/2)\/8)i)^1\/2"

"y2=(-(15\/8-(5959^(1\/2)\/8)i)^1\/2+(15\/8+(5959^(1\/2)\/8)i)^1\/2)\/2+3^(1\/2)i*((15\/8-(5959^(1\/2)\/8)i)^1\/2+(15\/8+(5959^(1\/2)\/8)i)^1\/2)\/2"

y3=(-(15/8-(5959^(1/2)/8)i)^1/2+(15/8+(5959^(1/2)/8)i)^1/2)/2-3^(1/2)i*((15/8-(5959^(1/2)/8)i)^1/2+(15/8+(5959^(1/2)/8)i)^1/2)/2

Answer x1,2,3=y1,2,3-1/2